OpenStudy (anonymous):

use the elimination method to solve the system of equations. 8x-4y=-2 8x+5y=7

OpenStudy (anonymous):

@Jamierox4ev3r

HanAkoSolo (jamierox4ev3r):

just turn everything in the first equation negative 8x-4y=-2 change that into..... -8x+4y=2

OpenStudy (anonymous):

9y=9 @Jamierox4ev3r

HanAkoSolo (jamierox4ev3r):

then, if you add -8x+4y=2 and 8x+5y=7 together, the 8x and -8x will cancel each other out and you are left with....nothing O.O

HanAkoSolo (jamierox4ev3r):

lol just kidding thats correct @Avon XD

OpenStudy (anonymous):

so the sollution set is 0,0 or ???

OpenStudy (anonymous):

@Jamierox4ev3r

HanAkoSolo (jamierox4ev3r):

lol no y=1, just as u have said. Now we need to solve for x

OpenStudy (anonymous):

ok so y=1

HanAkoSolo (jamierox4ev3r):

8x-4(1)=-2 8x+5(1)=7 8x-4=-2 8x+5=7 +4 +4 -5 -5 -------- ---------- 8x=2 8x=2 /8 /8 /8 /8 --- ----- x=1/4 x=1/4

HanAkoSolo (jamierox4ev3r):

so if x=\(\large\frac{1}{4}\) and y=1, then your final answer is (\(\large\frac{1}{4}, 1\))

HanAkoSolo (jamierox4ev3r):

Does that make sense @Avon ? :)

OpenStudy (anonymous):

nope i got lose but i c a lil bit

HanAkoSolo (jamierox4ev3r):

so we solved for y, then i used the y value (1) in both of the given equations in order to calculate the x value. Both of them equal \(\large\frac{1}{4}\) so that means that x=\(\large\frac{1}{4}\)